RC Meeting and Run Plan, Tuesday, February 25th 2020
- 1 RC Meeting Information
- 2 General updates
- 3 Current Run Plan
- 4 Notices in Effect for Shift Crew
- 5 Planned Accelerator Configuration Changes
- 6 Useful Links
RC Meeting Information
RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room
Tuesday 25th Feb. 2020.
- Beam last night and overnight was especially good, with fewer trips and 1/2 h periods without any trips!
- Last night lost raster IOC, had trouble rebooting it, then it came up with 0.5 Hz params instead of 5 Hz params. Resolved following procedure described here: https://logbooks.jlab.org/entry/3791260 with help from Mark Jones. Followup here: https://logbooks.jlab.org/entry/3791757
- Incredibly long time (1:20) to get into Hall C today, see https://logbooks.jlab.org/entry/3791678
- We'll try to squeeze in a 1 hour run with target out after the access today, to have a measure of the background just from the beryllium windows.
- Beam off at 8:30 for injector PSS work for 4 h, Hall C access should start at 9:00 AM. Actually started 10:15 because of crazy slow transition to controlled access.
- The helium leak was in Hall B so cryo will not sniff for it in Hall C today
- The techs (4-5 keys) have 3 pumps lined up to pump the xfer line IV on the HMS & SHMS, as well as the Moller magnet IV.
- the target experts are working on the target (laser access mode) at some point to investigate the polarization drops. In Laser Access at ~ 11 AM
- we'll pre-cool the Moller lines overnight Tuesday
- Cool-down the Moller Wednesday morning, or if cryo is part of the safety standown, Wed afternoon. We have coordinated with them on this already.
- Do the Moller measurement Thursday day
- Three hour-long safety stand-downown for accelerator Wednesday morning 9-12 AM (recent rumors suggest it may include engineering too)
- No access to the hall unless we were to go to restricted access much earlier than 9 and deal with the baggage to return to beam permit afterwards. We do not plan to do that. We will stay in power permit along with the accelerator. See https://logbooks.jlab.org/entry/3791697
- 3 hour stand-down Wed AM.
- Small changes in beam position have been deliberately made as well while we are at 30 uA on 3He (see https://logbooks.jlab.org/entry/3788864)
- Harp travel problem can't be fixed without taking the harps down (out). Limit switch is already at end of clearance. Way bigger job than we thought (so punt).
- In the meantime, since we take our harp scans with tune beam, we have been asking to deliberately move the beam 1-2 mm (1 mm is just barely enough) for the harp scan, then to move it back. We only care about the widths so this seems like a reasonable work-around for now. I modified the short beam-checkout wiki accordingly.
Not sure if it is a problem
- Raster energy nuisance alarms- fixed
- Brad added the Orbit lock indicator to our Hall C wall screen, making it easier to check if it's on or not
- Friday night adjustments by the crew chief (see above) may have ended the larger >1mm vertical excursions we have been having. But the jury's still out.
- Lots of noise on our energy since Hall A stole our 2A lock, and have been trying to get FFB turned on
- Since these are 2-3 MeV, ie dp/p ~ 2e-4 it's within specs and shouldn't impact our physics. So I haven't been complaining about it.
Opportunistic Problems to get done at some point
- Note there may need to be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames said at Wednesday's 1:30 meeting. See https://logbooks.jlab.org/entry/3788008
- The injector group may just move the spot- it remains to be seen.
- This could be checked when we go down for 4 h Tuesday to fix a PSS problem with the lasers in the injector
- Our QE is still reasonable, ~0.29 Monday AM (see https://logbooks.jlab.org/entry/3790779), was 0.32 a few days ago though
Opportunistic Access Jobs
- See above
Recent polarization measurements
- longitudinal polarization measurements after yesterday's target work in the hall and QWP angle optimization
- Tue 13:15: 47% https://logbooks.jlab.org/entry/3786686
- Tue 18:42: 52.2% https://logbooks.jlab.org/entry/3786898
- Tue 23:47: 54.5% https://logbooks.jlab.org/entry/3787175
- Wed 4:38: 55.0% https://logbooks.jlab.org/entry/3787175
- Wed 8:40: 54.9% https://logbooks.jlab.org/entry/3787285
- Wed 13:08 54.8% https://logbooks.jlab.org/entry/3787420
- Wed 18:15 54.8% https://logbooks.jlab.org/entry/3787605
- Wed 11 PM 55.3% https://logbooks.jlab.org/entry/3787781
- Wed midnight 51.9% https://logbooks.jlab.org/entry/3787926
- Thu 05:15 53.2% https://logbooks.jlab.org/entry/3787904
- Thu 10:34 53.6% https://logbooks.jlab.org/entry/3788085
- Thu 13:47 53.0% https://logbooks.jlab.org/entry/3788210
- Thu 21:00 56.0% https://logbooks.jlab.org/entry/3788428
- Fri 4:07 55.2% L https://logbooks.jlab.org/entry/3788613
- Fri 09:03 53.8% L https://logbooks.jlab.org/entry/3788795
- Fri 2:22 PM 52.5% L https://logbooks.jlab.org/entry/3788938
- Fri 19:50 51.9% L https://logbooks.jlab.org/entry/3789160
- Sat 01:15 51.6% L https://logbooks.jlab.org/entry/3789331
- Sat 06:32 53% L https://logbooks.jlab.org/entry/3789442
- Sat 11:50 54.9% L https://logbooks.jlab.org/entry/3789583
- Sat 19:31 43.2% L https://logbooks.jlab.org/entry/3789857
- Sat 23:08 50.2% L https://logbooks.jlab.org/entry/3789959
- Sun 04:06 55.3% L https://logbooks.jlab.org/entry/3790104
- Sun 09:21 55.5% L https://logbooks.jlab.org/entry/3790240
- Sun 14:17 55.7% L https://logbooks.jlab.org/entry/3790380
- Sun 19:03 55.7 L https://logbooks.jlab.org/entry/3790513
- Mon 00:25 55.2% L https://logbooks.jlab.org/entry/3790672
- Mon 05:47 49.6% L https://logbooks.jlab.org/entry/3790771
- Mon 10:59 50.8% L https://logbooks.jlab.org/entry/3790965
- Mon 15:51 50.4% L https://logbooks.jlab.org/entry/3791100
- Mon 21:22 50.9% L https://logbooks.jlab.org/entry/3791318
- Tue 02:59 50.7% L https://logbooks.jlab.org/entry/3791481
- Tue 8:30 50.4% L https://logbooks.jlab.org/entry/3791660
Current Run Plan
- Taking production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization until further notice. NMR every 5 hrs.
- Tuesday when we come out of the access, we'll take a 1h run with target out in order to measure the background from just the Beryllium windows.
- Next Moller measurement Thursday.
- If beam goes away for more than 2h, or if during the downtime, MCC does operations that may change then beam tuning (except work on transport magnets), then do beam recovery procedure (do YAG viewer part only if you are suspicious of the harp scan results)
- Check the position locks are on.
- Check CODA is running normally (had one strange hangup lately)
- CHeck online replays shortly after starting runs, & full replay results too.
- Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
- Always take 120Hz of EDTM
- Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! !! !! !! ATTN: SHIFT LEADERS !! !! Please follow the shift summary template on !! !! the laminated SL instructions. !! !! !! !! ATTN: TARGET OPERATORS !! !! All NMR measurement HCLOG entries should include: !! !! 1) A title that clearly lists polarization !! !! 2) Screenshots of both NMR_FieldSweep and !! !! the polarization fit !! !! !! !! For an example see the following entry: !! !! https://logbooks.jlab.org/entry/3768696 !! !! !! !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
- The transition to d2n will probably happen sometime the week of March 9-13.
- For the time being we will run exclusively with longitudinal polarization. Do NMR before and after transverse, and every 5 hours during longitudinal.
- Production at 30 degrees, HMS/SHMS=-3.5/-3.4 momentum.
When there is a down:
Notices in Effect for Shift Crew
- Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
- Magnet policies
- Spectrometer GUI (
go_magnets) should always be visible on Shift Leader computer, hcdesk3
- HMS Q1 must be ramped down in 50A steps in all cases
- Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
- If SHMS dipole trips again, take a screenshot of the PSU window before resetting the interlock and post it to hclog.
- Spectrometer GUI (
- Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
- The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
- Ramp rate lock instructions are here and added to checklist
- Temporary Laser instructions can be found here 
- Do not change SHMS momentum until we fixed the current lead valve
Planned Accelerator Configuration Changes
- Note there will likely be a recesiation of the photocathode in 10-14 d which could take 10-12 h, according to what Joe Grames reported at the 2/19 1:30 meeting.
- The injector group may just move the spot- it remains to be seen.
- This could be checked when we go down for 2-3 h in the next few days to fix a PSS problem with the lasers in the injector
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-188.8.131.52 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920